OpenStudy (cheesecakekitten):

@iGreen @AnswerMyQuestions @lordhelix8th What is 0.36 repeating expressed as the quotient of two integers in simplest form? Answer must be in the form of a fraction.

OpenStudy (rational):

what is repeating ? 6 or 36 ?

OpenStudy (cheesecakekitten):

36

OpenStudy (cheesecakekitten):

???

OpenStudy (phi):

what is 36.3636... - 0.3636... ?

OpenStudy (cheesecakekitten):

36

OpenStudy (phi):

ok, and (fixing a typo) I said x = 0.3636... and 100x = 36.3636 so 100 x - x = 36.36... - 0.36... or 100 x -x = 36 100 x's take away 1 x is 99x 99x = 36

OpenStudy (phi):

can you "solve for x" ?

OpenStudy (phi):

I would divide by 99 on both sides

OpenStudy (cheesecakekitten):

well @iGreen @AnswerMyQuestions?

OpenStudy (anonymous):

36 is the answer.

OpenStudy (igreen):

? Follow @phi 's instructions. They are excellent

OpenStudy (anonymous):

^^

OpenStudy (phi):

are you saying you can't divide by 99 on both sides? 99x = 36

OpenStudy (cheesecakekitten):

@sonic400539 hi, and the answer has to be a fraction. not a whole number

OpenStudy (cheesecakekitten):

i can i think

OpenStudy (igreen):

Divide 99 to both sides and keep it as a fraction, don't simplify to a decimal.

OpenStudy (cheesecakekitten):

ok

OpenStudy (cheesecakekitten):

ok, i can't do this. I got 1 / 363636

OpenStudy (cheesecakekitten):

ok so that equals 1x?

OpenStudy (phi):

\[ \frac{99x}{99} =\frac{36}{99} \]

OpenStudy (phi):

yes, 99/99 is 1 and 1x is just x

OpenStudy (phi):

remember we said x= 0.3636... way up top so if x = 0.3636... and x = 36/99 then 0.3636... = 36/99 notice we can simplify 36/99 (divide top and bottom by 9)

OpenStudy (kaley246):

Decide if the statement is always, sometimes, or never true. The graph of a quadratic equation is a straight line. (Points : 4) always sometimes never